# Searching for Structure: Investigating Emergent Communication with Large Language Models

### v2

@@ 5.3 Communication results in structure
Although the initial languages are unstructured, some form of structure emerges due to repeated learning and using (Figure 2). This mostly happens during the communication block where TopSim increases significantly across rounds () and Ngram decreases across rounds (). However, we also observe behaviour that is not human-like; the signals used to communicate become longer over the rounds (). This finding contradicts what is observed in human experiments, where we typically observe that messages become shorter and lie close to a theoretical frontier balancing expressivity and simplicity Piantadosi et al. (2011); Kirby et al. (2015). Finally, structured signals, i.e., signals with a higher degree of TopSim, benefit communicative success positively ().
These results extend the findings of Galke et al. (2023) that LLMs not only learn structured vocabularies better, but also naturally create languages that exhibit some form of structure when it is optimised for their inherent biases. In addition to LLMs’ struggle to learn impossible languages Kallini et al. (2024), the insight that reframing prompt instructions into a structured list improves the model response Mishra et al. (2022), and given that we do not impose pressure to induce structure, this surprising outcome may be the result of an apparent “structure bias” in LLMs.
@@ 5.4 Structure enables better generalisation
After the communication block, the agents engage in the final block of the simulation. Here they generate signals for all 27 stimuli using the vocabulary that has evolved after learning and communication. We find that high TopSim languages clearly allow for better generalisation (, Figure 3).
@@ 7 Discussion
Our findings show a mixed picture, agents comprised of LLMs can learn and use artificial languages in a referential game. They do so by optimising the initially holistic vocabulary to fit better with the preferences of their language model, resulting in increased regularity and structure (Table 1). Yet, we also observe some degeneracy, i.e. many-to-one mappings of signals and attributes, and non-humanlike behaviours such as a tendency to produce long signals. Iterated learning further increases the learnability of the vocabulary but also extrapolates these behaviours further. Despite not being able to directly compare our results to human data, these findings are loosely comparable to earlier work involving human participants Kirby et al. (2015); Raviv et al. (2019b) in which languages with similar properties emerge.

### v3

@@ 5.3 Communication results in structure
Although the initial languages are unstructured, some form of structure emerges due to repeated learning and use (Figure 2). This mostly happens during the communication block where TopSim increases significantly across rounds () and Ngram decreases across rounds (). This increase in structure benefits communicative success positively (). However, we also observe behaviour that is not human-like; the signals used to communicate become longer over the rounds (). This contradicts what is observed in human experiments, where we typically observe that messages become shorter and lie close to a theoretical frontier balancing expressivity and simplicity Piantadosi et al. (2011); Kirby et al. (2015).
These results extend the findings of Galke et al. (2023); LLMs not only learn structured vocabularies better but also naturally shape languages to have some form of structure when they are optimised for their inherent biases. As LLMs struggle to learn impossible languages Kallini et al. (2024), reframing prompt instructions into a structured list improves the model response Mishra et al. (2022), and given that we do not impose pressure to induce structure, the surprising outcome of our experiments may be the result of an apparent “structure bias” in LLMs.
@@ 5.4 Structure enables better generalisation
After the communication block, the agents engage in the final simulation block. Here they generate signals for all 27 stimuli using the vocabulary that has evolved after learning and communication. We find that high TopSim languages allow for better generalisation (, Figure 3).
@@ 7 Discussion
Our findings show a mixed picture, agents comprised of LLMs can learn and use artificial languages in a referential game. They do so by optimising the initially holistic vocabulary to fit better with the preferences of their language model, resulting in increased regularity and structure (Table 1). These human-like results are much in line with previous findings showing that structured languages can emerge from repeated interactions between interlocutors (i.a. Selten and Warglien, 2007; Verhoef et al., 2016; Nölle et al., 2018; Raviv et al., 2019a). Yet, we also observe some degeneracy, i.e. many-to-one mappings of signals and attributes, and non-humanlike behaviours such as a tendency to produce long signals. Iterated learning further increases the learnability of the vocabulary but also extrapolates these non-humanlike behaviours further. Despite not being able to directly compare our results to human data, these findings are loosely comparable to earlier work involving human participants Kirby et al. (2015); Raviv et al. (2019b) in which languages with similar properties emerge.